OpenStudy (anonymous):

Which ratios are proportional to 40 : 15? Choose all answers that are correct. A. 3 : 8 B. 8 : 3 C. 90 : 240 D. 120 : 45 Anyone help me, Please?~

OpenStudy (anonymous):

@pooja195 c:

pooja195 (pooja195):

whats the lowest 40 can be divided to?

OpenStudy (anonymous):

I think its B and D. But on and ill work that out c:

pooja195 (pooja195):

B is right

pooja195 (pooja195):

40/5=8 15/5=3

OpenStudy (caominhim):

the ratio will be proportional if you effect each number the same way ex: \[2:3 = 4:6\] because i just multiplied each side by 2

OpenStudy (anonymous):

:D

pooja195 (pooja195):

so your basically simplifying it :) you got this buddy! :D

OpenStudy (anonymous):

:0 ohkkay hold on ~.- (weird faces xD)

pooja195 (pooja195):

hmmm....you already have the answer though?

OpenStudy (anonymous):

I need a few more answers...... i dont know what they are but ill figure it out with some help i guess

pooja195 (pooja195):

sure ill help u out ^_^

OpenStudy (anonymous):

D?

pooja195 (pooja195):

wait you already have the answer LOL

pooja195 (pooja195):

look above

pooja195 (pooja195):

40/5=8 15/5=3

OpenStudy (anonymous):

the one he said wasnt an answer.....

pooja195 (pooja195):

my answer LOL you said i think B or D and i said B was correct...... 40/5=8 15/5=3

OpenStudy (anonymous):

B and?? o.O

pooja195 (pooja195):

LOL nvm xD hold on i feel dumb now yep D is also correct

OpenStudy (anonymous):

http://www.wolframalpha.com/ im borrowing this website for help LOL XD

pooja195 (pooja195):

LOL ok xD good luck! :)

OpenStudy (anonymous):

Alright welp cya later on XD